OpenStudy (anonymous):

int cosx tan (1+1/2sinx)dx

OpenStudy (lgbasallote):

cosx is a perfect differential of sinx....you can take that as a clue on how to manipulate it :D

OpenStudy (anonymous):

but I don't know how to do

OpenStudy (anonymous):

please help me

OpenStudy (dumbcow):

cos*tan = cos*sin/cos = sin -> sin(1+1/2sin) = sin +1/2sin^2

OpenStudy (anonymous):

thx a lot